OpenStudy (anonymous):

Solve this problem. A student needs at least 80% of all points on the tests in a class to get a B. There are three 100-point tests and a 250-point final. The student's test scores are 72%, 85% and 78% for the 100 point tests. What is the least score the student can make on the final and still get a B?

OpenStudy (anonymous):

Let x represent the % score on the last test. student needs at least 80% of all points on the tests in a class to get a B 235 + x = 320 x = 320-235 = 85 x = 85% Student needs an 85% on the last test 250 Question Test .85*250 = 212.5 Students needs a 'score' of at least 213 on last test

OpenStudy (anonymous):

not so sure if thats a correct answer:-)

OpenStudy (anonymous):

@ chris23 is correct. Student need 87% correct on the final.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@ jp-1203, don't you think you could spare a medal for chris23?

OpenStudy (anonymous):

hahahaha..thats okay rob